OpenStudy (anonymous):

point of intersection of 3x/4 +y/2 +3/4=0; x/4 -3y/2 +2/3=0

OpenStudy (anonymous):

Find the point of intersection of the two lines. \[\frac{ 3x }{ 4 }+ \frac{ y }{2}+ \frac{ 3 }{ 4}=0 ; \frac{ x }{ 4 }- \frac{ 3y }{2 }+\frac{ 2 }{ 3 }=0\]

OpenStudy (anonymous):

\[\frac{ 3x+2y+3 }{ 4}=\frac{ 3x-18y+8 }{12 }=0\] \[12(3x+2y+3)=4(3x-18y+8)=0\] \[36x+24y+36=12x-72y+32\] left side:\[36x+24y+36\]\[>>>>>\] 1)\[y=\frac{- 3 }{ 2 }(x+1)\] right side:\[12x-72y+32\]\[>>>>>>>>\]2)\[y=\frac{ 1 }{ 6 }x+\frac{ 4 }{9 }\] so,\[\frac{ 1 }{ 6 }x+\frac{ 4 }{9 }=\frac{- 3 }{ 2 }(x+1)\] solving for x: \[x=\frac{ -7 }{ 6 }\] substitute in 1)equation:\[y=\frac{ -3 }{ 2 }(\frac{ -7 }{6}+1)=\frac{ 1 }{ 4}\] so, intersection point :\[(\frac{ -7 }{ 6 },\frac{ 1 }{ 4 })\]